OpenStudy (anonymous):

solve by the elimination method 4x-9y=6.5 7y-2x=-4.5

OpenStudy (anonymous):

4x-9y=6.5 -2x+7y=-4.5 X 2 => 4x- 9y = 6.5 -4x+14y = 9 ---------------- 5y = 15.5 y = 3.1 => 4x-9*3.1=6.5 => 4x-27.9 = 6.5 => 4x = 34.4 => x = 8.6 Ans: x=8.6, y=3.1

OpenStudy (anonymous):

that solution is not correct

OpenStudy (anonymous):

Oh, sign mistake! I'll do it again :)

OpenStudy (anonymous):

4x-9y=6.5 -2x+7y=-4.5 X 2 => 4x- 9y = 6.5 -4x+14y = -9 ---------------- 5y = -2.5 y = -0.5 => 4x-9*-0.5=6.5 => 4x+4.5 = 6.5 => 4x = 2 => x = 0.5 Ans: x=0.5, y=-0.5

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

yes thank u.. can u help me with some others??